### v2

# Kreyòl-MT: Building MT for Latin American, Caribbean and Colonial African Creole Languages

Abstract
[…]
Kreyòl-MT: Building MT for Latin American, Caribbean and Colonial African Creole Languages
Nathaniel R. Robinson1 Raj Dabre3 Ammon Shurtz2 Rasul Dent4 Onenamiyi Onesi5 Claire Bizon Monroc4 Loïc Grobol6 Hasan Muhammad1 Ashi Garg1 Naome A. Etori7 Vijay Murari Tiyyala1 Olanrewaju Samuel8 Matthew Dean Stutzman2 Bismarck Bamfo Odoom1 Sanjeev Khudanpur1 Stephen D. Richardson2 Kenton Murray1 1Johns Hopkins University, USA; 2Brigham Young University, USA; 3National Institute of Information and Communications Technology, Japan; 4Inria Paris; 5Nile University of Nigeria; 6Université Paris Nanterre; 7University of Minnesota - Twin Cities, USA; 8University of Toronto nrobin38@jhu.edu
1 Motivation
From northern Brazil to the Gulf of Mexico, spanning an area including the Caribbean and Central American west coast, lies the Creole "civilizational region" Glissant (2008). One of its chief characteristics: a multiplicity of Creole languages, born from contact of African language speakers with European languages in the colonial era. Low-resource Creole languages are widely spoken here and throughout the world (Rickford and McWhorter, 2017; Mufwene, 2008; Bartens, 2021; Velupillai, 2015). Historic linguistic marginalization has stymied their technological advancement: few language technologies exist for these languages despite their many speakers (Lent et al., 2023).
Better MT could greatly benefit Creole language speakers. Many live in areas where their language is in the minority. Panama and Costa Rica are home to communities of West Indian descent who have maintained Creole languages (Conniff, 1983; Herzfeld, 1980). Large Haitian-speaking communities live in the Dominican Republic Zhong (2023), Chile, Mexico (Audebert, 2017), Brazil Terry and Mayes (2019), and the Bahamas (Perry, 2023; McCartney, 2013; Knowles, 2018). Language is one of the first obstacles to immigrants’ social integration, and many report daily reliance on MT Neto et al. (2020). The lack or low accuracy of such technologies can thus contribute to social exclusion.
Multiple Creole-speaking communities regularly fall victim to natural disasters Heinzelman and Waters (2010); Margesson and Taft-Morales (2010); Rasmussen et al. (2015); Look et al. (2019). As the frequency of Atlantic hurricanes may be accelerated by global climate change Hosseini et al. (2018), machine translation can provide useful tools to facilitate communication during international relief efforts (Lewis, 2010; Hunt et al., 2019).
Yet colonial-era stigmas dismissing Creole languages as broken or incomplete persist, and serve as justifications to advantage European languages at their expense (Alleyne, 1971; DeGraff, 2003). Association with lower economic status and limited use in official settings then inhibit data collection and Natural Language Processing (NLP) development for these languages (Lent et al., 2023). We expand on Lent et al. (2023, 2022) to unify community efforts in advancing Creole NLP.
In addition to meeting community needs, Creole language MT presents avenues of exploration for low-resource NLP. Many Creole languages have documented linguistic relationships with high-resource languages, as well as lexical and morphosyntactic proximity to each other (Rickford and McWhorter, 2017). (See § 3.) As such, they have potential for cross-lingual transfer Lent et al. (2023), a powerful technique for low-resource NLP Pfeiffer et al. (2020); Kim et al. (2019). This potential presents an opportunity to develop technologies for many Creole languages at once. But since state-of-the-art NLP methods rely on machine learning, this development is not possible without data. We present an expansive dataset for MT of Creole languages, as a meaningful first step towards developing their technologies. We contribute:
The largest, most genre-diverse MT datasets ever compiled for 41 Creole languages, including the first ever for 21 Creole languages
A public dataset of 11.6M aligned sentences and 3.4M monolingual sentences for 40 Creole languages111Visit our repository https://github.com/JHU-CLSP/Kreyol-MT for software, models, and data download.
Public models achieving state-of-the-art performance on a published Creole language benchmark for 26 language directions
2 Background and Related Work
Despite its potential benefits, previous Creole language MT development has been scarce. Google Translate,222https://translate.google.com a common MT interface, only supports one Latin American Creole language (Haitian). NLLB-200 (NLLB Team et al., 2022), a state-of-the-art MT model in the number of languages it supports (204), only supports five Creole languages. The emergence of large language models like ChatGPT,333chat.openai.com trained on massive unlabeled datasets, presents encouraging potential for more universal MT support. However, recent studies indicate that LLMs are unable to compete with supervised MT for low-resource languages (Robinson et al., 2023a; Zhu et al., 2023). Researchers have also speculated that LLMs’ MT capabilities in high-resource languages are due to curated bitexts in their training data (Briakou et al., 2022). Hence, developing bitext corpora for low-resource languages is still important in building MT for them.
Some prior works have approached Creole language MT, like the inclusion of early Haitian-English MT in the DIPLOMAT system (Frederking et al., 1997) and the 2011 Workshop on Statistical Machine Translation (WMT) (Callison-Burch et al., 2011). Creole languages that are the focus of more recent MT research include Haitian and Jamaican (Robinson et al., 2022a, 2023b), Naija (Adelani et al., 2022a; Ogueji and Ahia, 2019), Mauritian (Dabre and Sukhoo, 2022a), Sranan Tongo (Zwennicker and Stap, 2022), and Singlish (Wang et al., 2017; Liu et al., 2022). Additional NLP research has been conducted in sentiment analysis and named entity recognition for Naija (Oyewusi et al., 2020; Muhammad et al., 2022, 2023; Adelani et al., 2021); syntactic analysis for Singlish (Wang et al., 2017), Naija (Caron et al., 2019), and Martinican (Mompelat et al., 2022); and inference for Jamaican (Armstrong et al., 2022).
These works, while valuable steps forward, have been limited in their scope of languages. The prior work most comparable to ours is the recent CreoleVal (Lent et al., 2023), which focused on building datasets for 28 Creole languages. They focused on machine comprehension, relation classification, and MT. Lent et al.’s (2023) MT dataset is a significant contribution to our work. They do not publicly release their MT datasets, however. While portions of the data we collected (including part of the CreoleVal set) are not publicly releasable, we release 11.6M aligned bitext sentences and 3.4M monolingual sentences in 40 languages. The genre coverage of CreoleVal’s solely religious and Bible text data is another limitation shared by prior works, which can preclude general-purpose MT. For example, Robinson et al.’s (2023b) Haitian model trained on primarily religious texts achieves an impressive BLEU score of 68.0 on its same-genre test set, but a score of 14.7 on a Wikipedia-style test set. This indicates that Creole language MT models trained on specific genres may not be applicable to general domains. We provide the most diverse Creole language MT data yet, both in terms of languages and genres. (See § 3.3.)
Other multilingual MT works include some Creole languages. In addition to NLLB Team et al.’s (2022) five included Creole languages, Yuan et al.’s (2023a) multilingual Lego-MT model supports 8 Creole languages (out of 433 total). Our work is a significant expansion on this, making a meaningful contribution to the broader effort of low-resource dataset curation—including the WMT’23 shared task on the topic Sloto et al. (2023), which our work follows by involving noisy data filtering Minh-Cong et al. (2023) and document alignment Steingrímsson (2023).
3 Methodology and Dataset
The languages we include in our study are in Table 1. We retrieve each language’s vitality, official recognized status by a political entity (Off.), and number of speakers from Ethnologue.444https://www.ethnologue.com We label a language as a majority language (Maj.) if it is spoken by a majority of the population in one of its native countries or territories. This project includes languages from Latin American and Caribbean countries and territories. (See Figure 1.)
[…]
| Language | ISO | Glottocode | Native to… | Vitality | L1 | L2 | Off. | Maj. |
| Saint Lucian Patois | acf | sain1246 | Saint Lucia | Stable | - | ✗ | ✓ | |
| Bahamian Creole | bah | baha1260 | Bahamas | Stable | - | ✓ | ✓ | |
| Berbice Dutch | brc | berb1259 | Guyana | Extinct | ✗ | ✗ | ||
| Belizean Kriol | bzj | beli1260 | Belize | Institutional | - | ✗ | ✓ | |
| Miskito Coast Creole | bzk | nica1252 | Nicaragua | Institutional | - | ✗ | ✗ | |
| Garifuna | cab | gari1256 | Central America | Endangered | - | ✗ | ✗ | |
| Negerhollands | dcr | nege1244 | U.S. Virgin Islands | Extinct | ✗ | ✗ | ||
| Ndyuka | djk | ndyu1242 | Suriname, French Guiana | Stable | - | ✗ | ✗ | |
| Guadeloupean Creole | gcf | guad1243 | Guadeloupe | Stable | - | ✗ | ✓ | |
| Martinican Creole | gcf | mart1259 | Martinique | Stable | - | ✗ | ✓ | |
| French Guianese Creole | gcr | guia1246 | French Guiana | Stable | - | ✗ | ✓ | |
| Gullah | gul | gull1241 | South Carolina, Georgia | Endangered | - | ✗ | ✗ | |
| Creolese | gyn | creo1235 | Guyana | Stable | - | ✓ | ✓ | |
| Haitian | hat | hait1244 | Haiti | Institutional | ✓ | ✓ | ||
| San Andrés-Providencia | icr | sana1297 | Colombia | Stable | - | ✗ | ✗ | |
| Jamaican Patois | jam | jama1262 | Jamaica | Stable | ✗ | ✓ | ||
| Karipúna | kmv | kari1301 | Brazil | Endangered | - | ✗ | ✗ | |
| Louisiana Creole | lou | loui1240 | Louisiana | Endangered | - | ✗ | ✗ | |
| Media Lengua | mue | medi1245 | Ecuador | Endangered | - | ✗ | ✗ | |
| Papiamento | pap | papi1253 | Aruba, Curaçao, Bonaire | Institutional | ✓ | ✓ | ||
| San Miguel Creole | scf | sanm1305 | Panama | Extinct | ✗ | ✗ | ||
| Saramaccan | srm | sara1340 | Suriname, French Guiana | Stable | - | ✗ | ✗ | |
| Sranan Tongo | srn | sran1240 | Suriname | Institutional | ✗ | ✓ | ||
| Vincentian Creole | svc | vinc1243 | Saint Vincent | Stable | - | ✗ | ✓ | |
| Trinidadian Creole | trf | trin1276 | Trinidad | Stable | - | ✗ | ✓ | |
| Angolar | aoa | ango1258 | São Tomé and Príncipe | Endangered | - | ✗ | ✗ | |
| Saotomense | cri | saot1239 | São Tomé and Príncipe | Endangered | - | ✗ | ✗ | |
| Seychellois Creole | crs | sese1246 | Seychelles | Institutional | - | ✓ | ✓ | |
| Annobonese | fab | fada1250 | Equatorial Guinea | Stable | - | ✗ | ✗ | |
| Fanakalo | fng | fana1235 | South Africa | Endangered | 5.1 | ✗ | ✗ | |
| Pichi | fpe | fern1234 | Equatorial Guinea | Institutional | ✗ | ✗ | ||
| Ghanaian Pidgin | gpe | ghan1244 | Ghana | Institutional | ✗ | ✗ | ||
| Kabuverdianu | kea | kabu1256 | Cape Verde | Institutional | ✗ | ✓ | ||
| Krio | kri | krio1253 | Sierra Leonne | Institutional | ✗ | ✓ | ||
| Kituba | ktu | kitu1246 | Central Africa | Institutional | ✓ | ✓ | ||
| Mauritian | mfe | mori1278 | Mauritius | Institutional | ✗ | ✓ | ||
| Naija | pcm | nige1257 | Nigeria | Institutional | ✗ | ✓ | ||
| Guinea-Bissau Creole | pov | uppe1455 | Guinea-Bissau | Institutional | ✗ | ✓ | ||
| Principense | pre | prin1242 | São Tomé and Príncipe | Endangered | - | ✗ | ✗ | |
| Réunion Creole | rcf | reun1238 | Réunion | Stable | - | ✗ | ✓ | |
| Sango | sag | sang1328 | Central African Rep. | Institutional | ✓ | ✓ | ||
| Tok Pisin | tpi | tokp1240 | Papua New Guinea | Institutional | ✓ | ✓ | ||
| Cameroonian Pidgin | wes | came1254 | Cameroon | Institutional | - | ✗ | ✓ | |
As mentioned in § 2, Lent et al.’s (2023) CreoleVal work is most comparable to our own. But the set of languages we include differs from theirs. Their focus was on Creole languages in general, while ours is narrowed to those of the Americas, allowing us to dive deeper and include more languages with greater linguistic commonalities. The languages we include have shared patterns of historical formation. They generally have morphosyntactic commonalities with Niger-Congo languages (Kouwenberg and Lacharité, 2004; Castillo and Faraclas, 2006) and large-scale lexical overlap with Romance and Germanic languages (Valdman, 2000; Winford, 1997). For example, Haitian and Jamaican have extensively borrowed vocabulary from older forms of French and English, respectively, but they are morphosyntactically closer to Gbe, Kwa, and Igboid languages (Lefebvre, 2011; Brousseau, 2011; Seguin, 2020; Mufwene, 2002; Kouwenberg, 2008; Farquharson, 2012). Because we restrict our focus based on these linguistic traits, we also include some African Creole languages that are close phylogenetic relatives to our American focus languages, with some linguists arguing for a common ancestor McWhorter (2000).555We also include three distinct control languages: Sango and Kituba (African Creole languages with morphosyntactic proximity to Niger-Congo languages but no lexical proximity to European languages), and Tok Pisin (an Oceanic Creole language with lexical proximity to a Germanic language—English—but no relation to Niger-Congo languages.) For instance, Sierra Leonne’s Krio is likely related to Maroon Creole languages like Ndyuka and Saramaccan (Bhatt and Plag, 2012). Linguists have long noted the linguistic proximity of Louisiana Creole and French Guianese Creole to Creole languages of the Indian ocean (Mauritian, Seychellois, and Réunion Creole) Papen (1978). Papiamento is considered a descendant of Kabuverdianu by some Romero (2010), and Jamaican has the same phylogenetic relatives as Ghanaian Pidgin Amoako (1992); Cassidy (1966).
3.1 Collection methods
| Bitext | Monolingual | ||||||
| Web | |||||||
| Prev. pub. | aligned | articles | aligned | other | Prev. pub. | Web | |
| 14196475 | 21963 | 216756 | 18614 | 4683 | 2767602 | 607657 | 27081 |
We divide our dataset collection methodology into two stages: gathering and extraction. Creole languages were selected for an active gathering effort (underlined in Table 1). We excluded Haitian because its data were abundant in already identified sources. Further data were found for other languages, also included in Table 1. For each of the gathering languages, we performed the following steps. First, we searched research databases using query templates to track down already curated datasets. We searched “[language name]” on the ACL Anthology,666https://aclanthology.org followed by “[language name] machine translation”, “[language name] NLP”, and “[language name] translation” on Google Scholar.777https://scholar.google.com Query results from these search engines were typically prohibitively many, so we browsed top results until it became clear that the remainder were no longer relevant. (Individual data gatherers used best judgment for each language.) For some languages we multiplied queries to accommodate for alternate language names. (See Table 8 in Appendix A.) Second, for each language, we checked each of the following databases for parallel or monolingual corpora: OPUS (Tiedemann, 2012), Oscar (Ortiz Suárez et al., 2019) and LDC.888https://www.ldc.upenn.edu Third, we scoured the web for books with translations or additional resources. Fourth, we contacted researchers in the languages’ speaking communities for leads to potential data sources. Though we attempted various methods of contact whenever possible, response rates were low for this step. In all we gathered 107 sources such as anthology websites promoting cultural heritage or language revitalization, educational materials, and government documents.
After completing gathering, we moved to extraction. At this stage, we divided each of the gathered resources into groups, based on the data format. The six groups were: (1) parallel data previously published as a bitext, (2) web sources with aligned parallel sentences, (3) web sources containing unsegmented articles of text with translations, (4) PDF sources with aligned parallel sentences, (5) other PDF sources, and (6) sources of monolingual data. The amount of segments for each of these groups is summarized in Table 2, with a breakdown per language in Table 7 (Appendix A).
We immediately consolidated the previously published bitexts, including single-language resources from LAFAND-MT (Adelani et al., 2022a), KreolMorisienMT (Dabre and Sukhoo, 2022a) and the Caribe Trinidadian-English dataset (Smith, 2022). We gathered monolingual data from the MADLAD-400 clean corpus (Kudugunta et al., 2024) and JamPatoisNLI (Armstrong et al., 2022). To create novel bitexts, we then used the BeautifulSoup999https://www.crummy.com/software/BeautifulSoup/ and Selenium101010https://www.selenium.dev Python packages to extract text from web sources and the PyPDF2111111https://https://github.com/py-pdf/pypdf package for PDF sources. Organizing bitexts from sources with aligned parallel sentences was generally straightforward. When only unsegmented articles with translations were available, we segmented and aligned text based on punctuation and then manually corrected errors. Manual correction was performed by data extractors with sufficient proficiency in the languages involved. Our diverse team has some proficient speakers in our target languages (one L1 Martinican Creole speaker, two L2 Haitian speakers, one L2 Guadeloupean Creole speaker, one L2 Louisiana Creole speaker, two L2 Naija speakers, and one L2 Ghanaian Pidgin speaker). Though this did not cover anywhere near all of our included languages, proximity between Creole languages and related languages made the task doable.121212For instance, one Louisiana Creole speaker and one Haitian speaker—both also proficient in English, French, and linguistics—were able to manually correct alignment for Seychellois-English bitexts, due to Seychellois’ lexical overlap with Haitian, Louisiana Creole, and French. In general, the data extraction details varied for each source, since PDFs and websites have a wide variety of individualized styles and formats. In all, we extracted data from 39 of the 107 sources from our gathering stage.131313The remainder contained data in less accessible formats that we deemed too tedious to include in this work. We intend our dataset to be a living resource and hope to continually add to it as we encounter and format more data for Creole languages. Attributions for all data sources are in Appendix D.
As an appendage to our methodology, we incorporate data from published multilingual resources, including all bitexts from CreoleVal (Lent et al., 2023), Lego-MT (Yuan et al., 2023b), FLORES-200 dev and NLLB train data (NLLB Team et al., 2022), and AfricaNLP’23 (Robinson et al., 2023b) for any of our focus languages.141414We ensure not to include any data labeled for testing in our own data for training or tuning. We retrieved Bible translations from JHU (McCarthy et al., 2020a) for 18 languages, with up to four unique translations for each. We selected the four fullest English and French Bibles with reasonably modern text style from the same corpus, and formed up to eight bitexts for each language. We also contribute 360K previously unreleased unique parallel sentences from the Church of Jesus Christ of Latter-day Saints for hat, pap, and tpi with English. This data comes from religious sources, including scripture, instruction, discourses, humanitarian resources, genealogy, and administrative documents. We scraped small parallel corpora in the educational domain from APICS (Michaelis et al., 2013) for languages. We aggregated further published bitexts by crawling the web via Python’s mtdata package, which points to data from multiple online sources. Last, we added monolingual Wikipedia dumps for jam, gcr, gpe, and srn.
3.2 Data amounts
| Max. | Ours | Max. | Ours | Max. | Ours | |||
| prev. | (pub. / all) | prev. | (pub. / all) | prev. | (pub. / all) | |||
| acf | 15989 | 4406 / 23916 | gcf | 96 | 6467 / 6467 | mue | - | 147 / 147 |
| aoa | - | 198 / 198 | gcr | - | 1433 / 1433 | pap | 4898029 | 4968965 / 5363394 |
| bah | - | 327 / 327 | gpe | - | 223 / 223 | pcm | 31128 | 8084 / 47455 |
| brc | - | 222 / 222 | gul | 7990 | 266 / 8831 | pov | - | 480 / 480 |
| bzj | 23406 | 229 / 31002 | gyn | - | 258 / 258 | pre | - | 243 / 243 |
| bzk | - | 391 / 391 | hat | 4256455 | 5715227 / 6023034 | rcf | - | 285 / 285 |
| cab | 20879 | - / 20879 | icr | 15702 | 317 / 16774 | sag | 262334 | 260560 / 535310 |
| cri | - | 306 / 306 | jam | 25206 | 434 / 28713 | srm | 42303 | 440 / 59053 |
| crs | 222613 | 3186 / 225875 | kea | 129449 | 132931 / 132931 | srn | 583830 | 6620 / 615010 |
| dcr | - | 189 / 189 | kri | 50438 | 185 / 66736 | svc | - | 321 / 321 |
| djk | 45361 | 15266 / 68833 | ktu | 7886 | 175 / 10737 | tpi | 424626 | 451758 / 925648 |
| fab | - | 204 / 204 | lou | - | 1860 / 1860 | trf | - | 1691 / 1691 |
| fng | - | 160 / 160 | mart1259 | - | 5153 / 5153 | wes | - | 223 / 223 |
| fpe | - | 259 / 259 | mfe | 191909 | 25633 / 233320 |
Table 3 compares our own data size with the largest previously collected dataset for each language. Ours is the first dataset for languages, and the first public dataset for more. (See Table 5 for a comparison with individual prior works.)
[…]
3.3 Diversity of genres
As mentioned in § 2, prior works in Creole language MT are severely restricted in terms of genre, prohibiting them from general-purpose MT. We mitigate this by including a diversity of genres. Figure 2 shows that more than half of languages (26/41) cover at least two genres.151515Table 6 in Appendix A contains the same information in colorblind-friendly numerical form. The "Other/Mix" genre dominates charts for Haitian, Papiamento, and Total, because of the large NLLB train sets (sets we do not even include in our model training, see § 4). We include a final chart indicating the total with "Other/Mix" excluded, showing no majority among other genres. We acknowledge that our dataset’s genre diversity can still be improved in future iterations, but we highlight the dramatic increase in diversity it offers, compared to previous work.
4 Modeling Experiments
We now describe the experimental setup to show the utility of our datasets by training MT systems. We train and release models on our full datasets, but we also train on only the data we publicly release, to show its stand-alone utility. We experiment with cleaning our train sets versus leaving them as-are, and in some experiments we fine-tune mBART (Liu et al., 2020) rather than training from scratch.
4.1 Dataset Preprocessing
[…]
Cleaning: Each dev and test set was cleaned according to the GILT Leaders Forum’s Best Practices in Translation Memory Management.171717We release cleaning software on our repository https://github.com/JHU-CLSP/Kreyol-MT. We removed segments that were: empty, containing unbalanced brackets, mostly non-alphabetic characters, containing a source the same as the target, fewer than words, and containing a higher number of characters than standard deviations above the mean for data of that language. We normalized escaped characters and entities, white spaces, quotation marks, and the overall character set for each language. We removed any spurious characters that do not contribute semantically or syntactically in a segment and remove duplicates for all segments after cleaning to ensure there is no development and test set contamination. We toggle whether we thus clean train sets in both public and all, leading to a total of four configurations: public, public-cleaned, all, all-cleaned. Note that the dev and test sets, and hence models and their results, between public and all are not comparable, but those among cleaned and non-cleaned versions are (since the former variable implies independently split test sets, but latter variable toggles only whether train data were cleaned).
4.2 Implementation and Training
[…]
5 Results and Discussion
Figure 3 shows the performance of all models, tested for all available language pairs. We use chrF Popović (2015) because previous studies Mathur et al. (2020); Rei et al. (2020) show that it correlates better with human judgments than BLEU Papineni et al. (2002), and because semantic metrics such as COMET Rei et al. (2020) and BLEURT Sellam et al. (2020) lack support for low-resource languages.181818Other recent low-resource MT works also prioritize chrF Robinson et al. (2023a); Dabre and Sukhoo (2022b). Because many MT researchers have a more intuitive grasp of BLEU than chrF, we also include BLEU scores in Figure 5 of Appendix C.
As expected, high-resource pairs exhibited better translation quality than their lower resource counterparts, with some deviations from this trend possibly due to inevitable non-uniformity in test set genre. Some possibly overfit pairs like Gullah-French scored chrF up to 97.2, and some zero-shot pairs like Papiamento-Spanish scored as low as 4.1.
Despite the generally lower performance on their languages, we highlight the potential value of our dataset’s smallest bitexts. As stated in § 1, Creole languages’ linguistic relationships open the possibility of powerful cross-lingual transfer. Previous studies Ernštreits et al. (2022); Dabre et al. (2020); Arivazhagan et al. (2019); NLLB Team et al. (2022) have shown that MT models trained on numerous languages can often be adapted to translate new languages with few examples. Our results corroborate this finding; despite having only 391 parallel sentences to split for Miskito Coast Creole (bzk), our models achieved BLEU and chrF up to 43.7 and 60.4, respectively, translating into English. Vincentian Creole (svc) performed even better: up to 55.5 BLEU and 66.9 chrF into English, despite only 321 parallel sentences total. (Such scores would likely not be attainable via traditional bilingual training on such small sets Arivazhagan et al. (2019).) Strikingly, we perceived even higher performance on zero-shot Martinican-to-English translation, with 69.9 BLEU and 84.9 chrF. We hope future studies will reveal more about our dataset’s potential for low-resource cross-lingual transfer, and how it interfaces with dataset genre and diversity. This is an important exploration, not only to engineer better low-resource language systems, but as a scientific inquiry with general implications for low-resource language technologies.
Impact of corpora cleaning: In Figure 3, regardless of the use of public or all data, models trained on cleaned data (triangles and stars) typically outperform their counterparts trained on non-cleaned data (circles and diamonds), for higher resource languages. Cleaning eliminates noise and reduces variability in the especially noisy Creole datasets, which helps translation quality. However, cleaning can hurt performance by reducing the already scarce data for the lower-resource, more noisy corpora (like Pichi-English with 259 sentence pairs).
Does fine-tuning help? From Figure 3, a comparison of scratch and mBART models trained on cleaned data (squares and x’s) reveals that mBART fine-tuning helps both public and all data training. This fine-tuning helps more for out-of-Creole directions, with an average improvement of 8.1 chrF over scratch models on all data (compared to 1.4 chrF for into-Creole). Scratch models converged after 500k training steps of 33k-token batches, and mBART fine-tuning needed 500k-600k steps of only 8.2k-token batches. While this implies that mBART is data efficient, mBART fine-tuning is slow due to training about 700% more parameters. In total, mBART fine-tuning needed up to a week whereas scratch training needed only up to two days on the same hardware.
Comparison on existing test sets: Figure 4 compares our best models against results reported by CreoleVal (Lent et al., 2023) on relevant language pairs from the CreoleVal test sets. Since CreoleVal data is not public, we consider it apt to evaluate our best model (mBART) fine-tuned on all cleaned data. Overall, our models surpass CreoleVal models with +6.4 average chrF XENG and +14.1 average chrF ENGX. Our dataset is much larger and more domain diverse than CreoleVal’s; hence our improved results on most language pairs show that increasing data tends to be beneficial even if it reduces the domain specificity. We note that on a few language directions (8 of 34), CreoleVal’s model still beat ours, indicating the possibility for negative interference from a more diverse train set in some instances. In Appendix C we provide some additional comparisons of bilingual versus multilingual fine-tuning and evaluations on other existing benchmarks: Lego-MT Yuan et al. (2023a) and KreolMorisienMT Dabre and Sukhoo (2022a).
6 Conclusion
In this work, we compile the most comprehensive dataset to date for MT of Creole languages in the Americas. By aggregating disparate previous works and incorporating new data sources via scraping the web and PDFs, we expand MT datasets to 21 new languages and produce the largest and most genre-diverse in 20 more. We release translation models in 172 language directions, with 26/32 beating state-of-the-art benchmark performance, as well as a public dataset with 11.6M aligned bitexts and 3.4M monolingual sentences.
A large multilingual bitext like ours has potential to build the best yet or first ever MT models for many languages, something we accomplished on a surface level in this work but hope future works will continue. The data present a number of other potential uses, including: (1) training language models for applications like spelling correction Abdulrahman and Hassani (2022); Etoori et al. (2018); Al-Jefri and Mahmoud (2013); (2) availing textual data for applications like speech recognition and speech translation, which can be vital to low-literacy communities Robinson et al. (2022b); Gao et al. (2021); Rossenbach et al. (2020); (3) potential to study cross-lingual transfer between Creole languages and their phylogenetic relatives in yet unseen depth Robinson et al. (2023b); (4) research on the effects of linguistic data augmentation for small MT datasets (focusing on Creole languages’ unique linguistic position); (5) development of MT-assisted documentation tools for those languages that are endangered Bird and Chiang (2012); (6) the introduction of a common repository where any researchers in Creole NLP can accumulate datasets together and advance in collaboration Lent et al. (2023); etc. We hope that this work will provide valuable translation technologies to communities that have been historically under-served and inspire community-oriented efforts to further expand work on these low-resource languages.
Acknowledgements
We thank Suzanna Sia, Chris Emezue, Heather Lent, David Mortensen, Oliver Mayeux, Michael Gisclair, Arya McCarthy, Ruth-Ann Armstrong, Carter Charles, Jeff Allen, Jamell Dacon, and Fritz-Carl Morlant for their contributions to the ideas and processes of this work.
Limitations
Across languages, writing systems change over time due to linguistic changes, such as the loss of distinctions between sounds; or metalinguistic changes, such as the desire to associate a speech community with a more prestigious one (or conversely show that the speech community is distinct). This concern is exacerbated for Creole languages, which tend to have very recent and often still developing standardization processes (Deuber and Hinrichs, 2007; Valdman, 2005; Rajah-Carrim, 2009). For several of our languages, especially Louisiana Creole and French Guianese Creole, we rely extensively on texts that were written more than 100 years ago and thus use spelling systems that have been partially or wholly superseded by new systems. In principle, it is possible to rewrite such texts with more recent conventions using language-specific scripts, but we opted to use the original orthographies for this work to keep the processing pipelines as similar as possible across languages. In the course of our data collection efforts, we identified several sources which we did not have time to process. In the future, we intend to continue adding new sources into our dataset, with a preference for those which are already publicly available.
Currently in our experiments, the newly created development and test scores are not comparable across the public and all splits due to their independent splitting processes (and hence dissimilar test sets). Our future work will focus on having development and test sets that are common across both splits, regardless of whether the training data in said splits are cleaned or not, for consistent comparisons.
Ethics Statement
Because Creole languages have frequently been the target of "othering" and marginalization (DeGraff, 2005; Lent et al., 2023), it is important to approach Creole language technologies with sensitivity. From a linguistic standpoint, the question of "What exactly makes a language Creole?" has contribued to the lack of prestige that many Creole languages currently face. For this study, we use existing literature to identify languages that have been considered Creole, but do not seek to assert a singular "Creole essence."
As with any other linguistic community, there are considerable differences in opinion concerning the desirability of MT in various Creole-speaking communities. We acknowledge that MT technologies do not inherently benefit all Creole language speakers. Many of them can already use existing MT tools in a different language, lessening the immediate benefit of MT tools in the relevant Creole. Others may be concerned about machine translation displacing human translators, or view the entire concept of MT as offensive, as it directly broaches the subject of linguistic differences between their languages and European languages (differences which are still broadly stigmatized). We also acknowledge that the intended use of some of the resources we collected may not have been MT. We do not wish to undermine the original purposes of anthology-style data but hope our work will support these endeavors.
We also acknowledge that the texts we have assembled, especially those which are older, religious and/or political, reflect many different viewpoints that may be considered dated, contested, or offensive in some cultural contexts. This is a natural part of the data collection process, but it is not an endorsement of the content of any given text. We did not seek to remove such viewpoints from our data, as they are culture-specific. Another risk of data collection is the inclusion of personally identifiable information that may pose a risk to some users. This is a particular problem with a commonly used Haitian-English bitext from WMT 2011 (Callison-Burch et al., 2011). Though it is difficult to avoid data contamination in this vein completely, we avoid including this dataset to mitigate this risk. We also acknowledge the potential for bias in our dataset, since it is not perfectly balanced in terms of genres and topics. We encourage more application-oriented work in the future to report MT results broken down by test set genre.
In our conception, the primary beneficiaries of Creole MT technologies would be monolingual Creole language speakers. Many monolingual Creole language speakers have limited literacy and would perhaps benefit from speech translation systems more than text-based systems. As such, we encourage future work in the area of speech technologies for Creole languages and hope the textual materials and models we provide in this work can be of use to that end.
Lastly, ChatGPT was used to assist software writing for this project. We acknowledge the ethical implications of LLM use are still being understood.
Appendix A Additional Data Information
Table 5 contains bitext sizes for individual previous publications for each language, compared to our own datasets (summarized in Table 3). Table 6 provides numerical values for our dataset’s genre composition (corresponding to Figure 2). Table 7 gives language-by-language details of the types of data we extracted in our collection methodology (summarized in Table 2). Table 8 shows alternate names for languages with more than one used in data gathering.
Appendix B Experimental Details
We describe some additional details related to training and evaluation.
B.1 mBART-50 Token Repurposing
mBART-50 has 51 special tokens corresponding to exactly 51 languages in our experiments, of which 41 are Creoles. Therefore, we simply repurpose these tokens where we fix the tokens for English, French, Spanish and Portuguese to the corresponding tokens in the mBART-50 tokenizer and then randomly assigned other tokens to Creoles. We found that this only affects initial training, as the model has to re-learn that the token is used to translate into another language.
B.2 Training and Convergence
[…]
B.3 Hyperparameters
| Optimizer | AdamW |
| Learning Rate | () |
| Weight Decay | |
| #encoder/decoder layers | () |
| hidden size | () |
| FFN size | () |
[…]
| data sampling temperature202020We selected 2.0 to oversample the smaller corpora only slightly. We noticed in our peliminary experiments that the standard temperature of 5.0 caused the higher resource pairs to suffer. | |
[…]
| dropout | () |
| label smoothing |
The relevant training hyperparameters are given in Table 4.
B.4 Additional Models Trained
We train a simpler version of m2m-mBART focusing only on Haitian–English bidirectional translation to determine whether its better to focus on fewer language pairs during fine-tuning. The key changes in hyperparameters are learning rate () and dropout ().
Appendix C Additional Results
C.1 BLEU Scores For Our Test Sets
Figure 5 shows the BLEU scores on our newly proposed test sets. These are analogous to the chrF scores presented in Figure 3.
C.2 Bilingual vs Multilingual mBART fine-tuning
Figure 6 shows the results of fine-tuning mBART-50 only on Haitian–English vs on multilingual data, for the public split mentioned in Section 4.1. We can see that bilingual models are inferior to multilingual models. The same trend exists for the all split.
C.3 Lego-MT and KreolMorisienMT Results
Figure 7 shows results of our models on the KreolMorisienMT Dabre and Sukhoo (2022a) test sets, and Figure 8 shows results of the same models on the Lego-MT Yuan et al. (2023a) test sets. Yuan et al. (2023a) did not report results on their own test splits; hence we do not compare our models’ performance directly with theirs. Once again, our fine-tuned models give the best performance.
| CreoleVal | JHU | Lego-MT | FLORES | AfricaNLP | NLLB | Ours | Ours | |
| Public? | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✓ | ✗ ✓ |
| acf | 7864 | 15989 | - | - | - | - | 4406 | 23916 |
| aoa | - | - | - | - | - | - | 198 | 198 |
| bah | - | - | - | - | - | - | 327 | 327 |
| brc | - | - | - | - | - | - | 222 | 222 |
| bzj | 14911 | 23406 | - | - | - | - | 229 | 31002 |
| bzk | - | - | - | - | - | - | 391 | 391 |
| cab | - | 20879 | - | - | - | - | - | 20879 |
| cri | - | - | - | - | - | - | 306 | 306 |
| crs | 222613 | 5055 | - | - | - | - | 3186 | 225875 |
| dcr | - | - | - | - | - | - | 189 | 189 |
| djk | 45361 | 23748 | 7868 | - | - | - | 15266 | 68833 |
| fab | - | - | - | - | - | - | 204 | 204 |
| fng | - | - | - | - | - | - | 160 | 160 |
| fpe | - | - | - | - | - | - | 259 | 259 |
| gcf | - | - | 96 | - | - | - | 6467 | 6467 |
| gcr | - | - | - | - | - | - | 1433 | 1433 |
| gpe | - | - | - | - | - | - | 223 | 223 |
| gul | 7870 | 7990 | - | - | - | - | 266 | 8831 |
| gyn | - | - | - | - | - | - | 258 | 258 |
| hat | 210593 | 72354 | 477048 | 2006 | 179435 | 4256455 | 5715227 | 6023034 |
| icr | 7799 | 15702 | - | - | - | - | 317 | 16774 |
| jam | 7988 | 25206 | 26 | - | 5118 | - | 434 | 28713 |
| kea | - | - | - | 2009 | - | 129449 | 132931 | 132931 |
| kri | 50438 | 23740 | - | - | - | - | 185 | 66736 |
| ktu | 7886 | 5055 | - | - | - | - | 175 | 10737 |
| lou | - | - | - | - | - | - | 1860 | 1860 |
| mart1259 | - | - | - | - | - | - | 5153 | 5153 |
| mfe | 191909 | 23625 | 399 | - | - | - | 25633 | 233320 |
| mue | - | - | - | - | - | - | 147 | 147 |
| pap | 397354 | 5018 | 269 | 2009 | - | 4898029 | 4968965 | 5363394 |
| pcm | 31128 | 15905 | - | - | - | - | 8084 | 47455 |
| pov | - | - | - | - | - | - | 480 | 480 |
| pre | - | - | - | - | - | - | 243 | 243 |
| rcf | - | - | - | - | - | - | 285 | 285 |
| sag | 262334 | 16952 | 9 | 2009 | - | 235749 | 260560 | 535310 |
| srm | 42303 | 23531 | - | - | - | - | 440 | 59053 |
| srn | 583830 | 24569 | - | - | - | - | 6620 | 615010 |
| svc | - | - | - | - | - | - | 321 | 321 |
| tpi | 398341 | 81595 | 70 | 2009 | - | 424626 | 451758 | 925648 |
| trf | - | - | - | - | - | - | 1691 | 1691 |
| Lang | Bible | Educational | Legal | Narrative | News | Religious | Wikipedia | Other/Mix |
| acf | 15989 | 4406 | 0 | 0 | 0 | 0 | 0 | 33778 |
| aoa | 0 | 198 | 0 | 0 | 0 | 0 | 0 | 0 |
| bah | 0 | 327 | 0 | 0 | 0 | 0 | 0 | 0 |
| brc | 0 | 222 | 0 | 0 | 0 | 0 | 0 | 0 |
| bzj | 54933 | 229 | 0 | 0 | 0 | 0 | 0 | 10213 |
| bzk | 0 | 391 | 0 | 0 | 0 | 0 | 0 | 0 |
| cab | 20879 | 0 | 0 | 0 | 0 | 0 | 0 | 47471 |
| cri | 0 | 306 | 0 | 0 | 0 | 0 | 0 | 0 |
| crs | 5055 | 273 | 443948 | 15719 | 4141 | 0 | 0 | 279331 |
| dcr | 0 | 189 | 0 | 0 | 0 | 0 | 0 | 0 |
| djk | 23815 | 7398 | 0 | 0 | 0 | 0 | 0 | 66491 |
| fab | 0 | 204 | 0 | 0 | 0 | 0 | 0 | 0 |
| fng | 0 | 160 | 0 | 0 | 0 | 0 | 0 | 0 |
| fpe | 0 | 259 | 0 | 0 | 0 | 0 | 0 | 0 |
| gcf | 1559 | 304 | 0 | 4446 | 0 | 0 | 0 | 158 |
| gcr | 879 | 159 | 0 | 2388 | 0 | 0 | 15141 | 0 |
| gpe | 0 | 223 | 0 | 0 | 0 | 0 | 12425 | 0 |
| gul | 7990 | 262 | 0 | 0 | 0 | 0 | 0 | 579 |
| gyn | 0 | 258 | 0 | 0 | 0 | 0 | 0 | 0 |
| hat | 71958 | 9359 | 0 | 4 | 0 | 115583 | 5452 | 5828317 |
| icr | 15702 | 317 | 0 | 0 | 0 | 0 | 0 | 755 |
| jam | 18420 | 233 | 0 | 0 | 0 | 0 | 4588 | 29647 |
| kea | 0 | 6108 | 0 | 0 | 229 | 0 | 0 | 132314 |
| kri | 16113 | 185 | 0 | 0 | 0 | 0 | 0 | 99454 |
| ktu | 5055 | 175 | 0 | 0 | 0 | 0 | 0 | 62584 |
| lou | 0 | 668 | 0 | 1192 | 0 | 0 | 0 | 0 |
| mart1259 | 0 | 283 | 0 | 0 | 0 | 0 | 0 | 4870 |
| mfe | 23624 | 258 | 0 | 274 | 0 | 0 | 0 | 277331 |
| mue | 0 | 147 | 0 | 0 | 0 | 0 | 0 | 0 |
| pap | 5018 | 2996 | 0 | 0 | 0 | 65573 | 92 | 7430445 |
| pcm | 15905 | 253 | 0 | 0 | 0 | 0 | 0 | 31297 |
| pov | 0 | 480 | 0 | 0 | 0 | 0 | 0 | 0 |
| pre | 0 | 243 | 0 | 0 | 0 | 0 | 0 | 0 |
| rcf | 0 | 285 | 0 | 0 | 83659 | 0 | 0 | 63975 |
| sag | 16952 | 192 | 0 | 0 | 0 | 0 | 0 | 518166 |
| srm | 23531 | 440 | 0 | 0 | 0 | 0 | 0 | 57013 |
| srn | 24567 | 6607 | 0 | 0 | 0 | 0 | 4600 | 766364 |
| svc | 0 | 321 | 0 | 0 | 0 | 0 | 0 | 0 |
| tpi | 81178 | 62 | 0 | 0 | 0 | 25018 | 7 | 819383 |
| trf | 0 | 174 | 0 | 0 | 0 | 0 | 0 | 1517 |
| wes | 0 | 223 | 0 | 0 | 0 | 0 | 0 | 0 |
| Total | 449122 | 45777 | 443948 | 24023 | 88029 | 206174 | 42305 | 16561453 |
| Bitext | Monolingual | |||||||
| Web | ||||||||
| Lang | Prev. pub. | aligned | articles | aligned | other | Prev. pub. | Web | |
| acf | 19510 | 0 | 0 | 4406 | 0 | 30257 | 0 | 0 |
| aoa | 0 | 198 | 0 | 0 | 0 | 0 | 0 | 0 |
| bah | 0 | 327 | 0 | 0 | 0 | 0 | 0 | 0 |
| brc | 0 | 222 | 0 | 0 | 0 | 0 | 0 | 0 |
| bzj | 30773 | 229 | 0 | 0 | 0 | 0 | 34373 | 0 |
| bzk | 0 | 391 | 0 | 0 | 0 | 0 | 0 | 0 |
| cab | 20879 | 0 | 0 | 0 | 0 | 47471 | 0 | 0 |
| cri | 0 | 306 | 0 | 0 | 0 | 0 | 0 | 0 |
| crs | 222690 | 273 | 0 | 0 | 2912 | 61696 | 445177 | 15719 |
| dcr | 0 | 189 | 0 | 0 | 0 | 0 | 0 | 0 |
| djk | 61435 | 7398 | 0 | 0 | 0 | 28871 | 0 | 0 |
| fab | 0 | 204 | 0 | 0 | 0 | 0 | 0 | 0 |
| fng | 0 | 160 | 0 | 0 | 0 | 0 | 0 | 0 |
| fpe | 0 | 259 | 0 | 0 | 0 | 0 | 0 | 0 |
| gcf | 64 | 4844 | 1559 | 0 | 0 | 0 | 0 | 0 |
| gcr | 0 | 159 | 880 | 0 | 394 | 0 | 15140 | 1994 |
| gpe | 0 | 223 | 0 | 0 | 0 | 0 | 12425 | 0 |
| gul | 8569 | 262 | 0 | 0 | 0 | 0 | 0 | 0 |
| gyn | 0 | 258 | 0 | 0 | 0 | 0 | 0 | 0 |
| hat | 5900275 | 165 | 122594 | 0 | 0 | 0 | 7639 | 0 |
| icr | 16457 | 317 | 0 | 0 | 0 | 0 | 0 | 0 |
| jam | 28311 | 233 | 169 | 0 | 0 | 19756 | 4419 | 0 |
| kea | 131454 | 484 | 860 | 0 | 133 | 0 | 229 | 5491 |
| kri | 66551 | 185 | 0 | 0 | 0 | 49016 | 0 | 0 |
| ktu | 10562 | 175 | 0 | 0 | 0 | 57077 | 0 | 0 |
| lou | 0 | 440 | 0 | 228 | 1192 | 0 | 0 | 0 |
| mart1259 | 0 | 231 | 0 | 4870 | 52 | 0 | 0 | 0 |
| mfe | 232788 | 258 | 0 | 274 | 0 | 68167 | 0 | 0 |
| mue | 0 | 147 | 0 | 0 | 0 | 0 | 0 | 0 |
| pap | 5294750 | 146 | 65665 | 2833 | 0 | 2140730 | 0 | 0 |
| pcm | 47202 | 253 | 0 | 0 | 0 | 0 | 0 | 0 |
| pov | 0 | 480 | 0 | 0 | 0 | 0 | 0 | 0 |
| pre | 0 | 243 | 0 | 0 | 0 | 0 | 0 | 0 |
| rcf | 0 | 285 | 0 | 0 | 0 | 60098 | 83659 | 3877 |
| sag | 535118 | 192 | 0 | 0 | 0 | 0 | 0 | 0 |
| srm | 58613 | 440 | 0 | 0 | 0 | 21931 | 0 | 0 |
| srn | 608397 | 606 | 4 | 6003 | 0 | 182532 | 4596 | 0 |
| svc | 0 | 321 | 0 | 0 | 0 | 0 | 0 | 0 |
| tpi | 900560 | 63 | 25025 | 0 | 0 | 0 | 0 | 0 |
| trf | 1517 | 174 | 0 | 0 | 0 | 0 | 0 | 0 |
| wes | 0 | 223 | 0 | 0 | 0 | 0 | 0 | 0 |
| Total | 14196475 | 21963 | 216756 | 18614 | 4683 | 2767602 | 607657 | 27081 |
| ISO | Name 1 | Name 2 | Name 3 | Name 4 |
| gcf | French Antillean | Guadeloupean Creole | Martinican | - |
| gcr | French Guianese | Kriyòl Gwiyannen | Kriyòl Lagwiyann | Gwiyannen |
| djk | Ndyuka | Eastern Maroon Creole | Aukan | Nengee |
| kmv | Karipúna | Amapá Creole | Uaçá Creole | - |
| bzk | Miskito Coast Creole | Nicaraguan Creole English | - | - |
| pcm | Naija | Nigerian Pidgin | - | - |
| kea | Cape Verdean | Kabuverdianu | - | - |
| mfe | Mauritian Creole | Morisyen | - | - |
| crs | Seychellois | Seselwa | Kreol Sesel | - |
Appendix D Attributions
We provide exact attributions for all our data sources here. We list them with string identifiers that we used to distinguish them in our own data organization. See our repository https://github.com/JHU-CLSP/Kreyol-MT for data downloading instructions.
D.1 Resources for Bitexts
APiCS - Atlas of Pidgin and Creole Language Structures Michaelis et al. (2013)
Languages: dcr, icr, bzj, pov, fng, trf, rcf, fpe, kri, pap, gcf, gpe, tpi, crs, ktu, pre, fab, bah, srm, gyn, djk, brc, sag, aoa, pcm, svc, mart1259, bzk, cri, gcr, kea, wes, hat, lou, srn, jam, mue, mfe, gul
Links: https://apics-online.info/
AfricaNLP-2023 - Aligned sentenced pairs from Robinson et al. (2023b)
Languages: hat, jam
bible_uedin - bible-uedin-v1, Parallel corpus created from translations of the Bible Christodouloupoulos and Steedman (2015)
Languages: djk
Bidze2019 - Seychelles Government Budget For the Fiscal Year 2019, Office of the President of The Republic of Seychelles
Languages: crs
Bidze2021 - Seychelles Government Budget For the Fiscal Year 2021, Office of the President of The Republic of Seychelles
Languages: crs
boston-food-forest - Boston Food Forest Coalition flyers translations
Languages: kea
CJCLDS - Online library of The Church of Jesus Christ of Latter-day Saints
Languages: pap, hat, tpi
Links: https://www.churchofjesuschrist.org/study?lang=pap (Link to full LDC dataset available on our repository: https://github.com/JHU-CLSP/Kreyol-MT.)
CREOLORAL - Martinican and Guadeloupean oral corpus with annotations Herby (2012)
Languages: gcf
Confiant-Dictionary - Dictionnaire Créole Martiniquais - Français, Raphaël Confiant Confiant (2007)
Languages: mart1259
CreoleVal Lent et al. (2023)
Languages: djk, kri, icr, pap, hat, bzj, sag, ktu, acf, srn, pcm, tpi, jam, crs, mfe, gul, srm
dicoNengee - Dictionnaire Nengee - Français - English
Languages: djk
FLORES-200 NLLB Team et al. (2022)
Languages: kea, pap, hat, sag, tpi
folklore - Excerpts from Le folklore de l’Ile-Maurice (texte créole et traduction française) Baissac (1888)
Languages: mfe
fortier - Excerpts from Louisiana Folk-tales: In French Dialect and English Translation Fortier (1895)
Languages: lou
GoiloText - Papiamentu Textbook, E.R. Goilo
Languages: pap
JHU - The Johns Hopkins University Bible Corpus McCarthy et al. (2020b)
Languages: djk, kri, icr, pap, hat, bzj, sag, cab, ktu, acf, srn, pcm, tpi, jam, crs, mfe, gul, srm
kapes - Corrections of the "Certificat d’aptitude au professorat de l’enseignement du second degré" (CAPES) exam for Martinican and Guadeloupean Creole
Languages: gcf, mart1259
KreolMorisienMT Dabre and Sukhoo (2022a)
Languages: mfe
LAFANDMT Adelani et al. (2022b)
Languages: pcm
LegoMT Yuan et al. (2023b)
Languages: djk, pap, gcf, hat, sag, tpi, jam, mfe
mindelo - Online dictionary
Languages: kea
MIT-Haiti, MIT-Haiti Initiative Lent et al. (2023)
Languages: hat
MiBelNouvel - Translation of the Gospel of John in Guadeloupean Creole
Languages: gcf
Languages: hat
Languages: hat, tpi, sag
PwovebKreyol - Proverbes & expressions créoles
Languages: gcf
QCRI - Abdelali et al. (2014)
Languages: pap
QED - Lamm et al. (2021)
Languages: mfe, hat, pap, sag
quentin - Excerpts from Introduction à l’histoire de Cayenne ; suivie d’un Recueil de contes, fables et chansons en créole de Saint-Quentin (1872)
Languages: gcr
SIL-Suriname - Languages of Suriname, SIL
Languages: djk, srm
Saint_Lucia_Ministry_of_Ed - Kwéyòl Dictionary, Ministry of Education Government of Saint Lucia
Languages: acf
TEC-English - Trinidad English Creole to English Dataset, University of the West Indies Smith (2022)
Languages: trf
TED2020 - TED and TED-X transcripts Reimers and Gurevych (2020)
Languages: hat
Tatoeba - Tatoeba database
Languages: pap, gcf, sag, srn, tpi, jam, mfe, hat
TiLiv - Ti Liv Kreyol, A learner’s guide to Louisiana creole Guillory-Chatman et al. (2020)
Languages: lou
Ubuntu - Ubuntu Translations
Languages: hat, pap
Links: https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en-ht.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_AU-ht.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_CA-ht.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_GB-ht.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en-pap.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_AU-pap.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_CA-pap.txt.zip, https://object.pouta.csc.fi/OPUS-Ubuntu/v14.10/moses/en_GB-pap.txt.zip
Wikimedia
Languages: pap, hat, tpi, jam, srn, gcr
Links: https://en.wikipedia.org/
Wikipedia
Languages: gcf
Wortubuku - Wortubuku fu Sranan Tongo, Sranan Tongo - English Dictionary Pinas et al. (2007)
Languages: srn
XLEnt El-Kishky et al. (2021)
Languages: hat
YouVersion–Bible - Life.Church online Bible
Languages: gcr
D.2 Resources for Monolingual Corpora
Anacao - Articles from the A Nação newspaper
Languages: kea
Links: https://www.anacao.cv/,
atipa - Excerpts from the book Atipa Parépou and Fauquenoy (1987)
Languages: gcr
Belizean - Life.Church online Bible
Languages: bzj
Creolica - Online corpus of Seychellois Créole
Languages: rcf, crs
Fonologia - Extracted interviews from Rodrigues (2007)
Languages: kea
graelo - Wikipedia dumps
Languages: gcr
JamPatoisNLI - Armstrong et al. (2022)
Languages: jam
KreolMorisienMT - Dabre and Sukhoo (2022b)
Languages: mfe
MADLAD-400 - Kudugunta et al. (2024)
Languages: mfe, pap, crs, kri, srm, jam, srn, djk, ktu, acf, rcf, cab, bzj
MIT-Haiti - Learning Resources from the MIT-Haiti initiative Lent et al. (2023)
Languages: hat
National - Official minutes of the Sittings of the House, The National Assembly of Seychelles
Languages: crs
Seychelles - Articles from the NATION newspaper, National Information Services Agency (NISA)
Languages: crs
Links: http://nation.sc/
temoignages - Articles from the Témoignages newpaper
Languages: rcf
Wikidumps
Languages: jam, gpe, gcr, srn
Wikimedia
Languages: srn